# Comparing Apples to Oranges: A Dataset & Analysis of LLM Humour Understanding from Traditional Puns to Topical Jokes

### v1

@@ Abstract
Humour, as complex language form, is derived from myriad aspects of life, whilst existing work on computational humour has focussed almost exclusively on short pun-based jokes. In this work, we investigate whether the ability of Large Language Models (LLMs) to explain humour depends on the particular humour form. We compare models on simple puns and more complex topical humour that requires knowledge of real-world entities and events. In doing so, we curate a dataset of 600 jokes split across 4 joke types and manually write high-quality explanations. These jokes include heterographic and homographic puns, contemporary internet humour, and topical jokes, where understanding relies on reasoning beyond ”common sense”, rooted instead in world knowledge regarding news events and pop culture. Using this dataset, we compare the zero-shot abilities of a range of LLMs to accurately and comprehensively explain jokes of different types, identifying key research gaps in the task of humour explanation. We find that none of the tested models (inc. reasoning models) are capable of reliably generating adequate explanations of all joke types, further highlighting the narrow focus of most works in computational humour on overly simple joke forms.111CONTENT WARNING: Some of the presented jokes may be considered offensive or distasteful to different individuals.
@@ A Dataset & Analysis of LLM Humour Understanding from Traditional Puns to Topical Jokes
Tyler Loakman1, William Thorne1, and Chenghua Lin2 1Department of Computer Science, University of Sheffield, UK 2Department of Computer Science, University of Manchester, UK {tcloakman1,wthorne1}@sheffield.ac.uk chenghua.lin@manchester.ac.uk
@@ 1 Introduction
The perception and understanding of humour is essential to human interaction and entertainment. Consequently, the ability to correctly process jokes is paramount to achieving a human-like understanding of ambiguity and social context in language. However, most existing work on computational humour have narrowly focussed on the simpler task of humour detection Loakman et al. (2023a); Meaney et al. (2021); Miller et al. (2017), which allows models to learn from confounding signals to detect certain joke types Lima Inácio et al. (2023). On the other hand, humour explanation tasks a computational model with the more challenging objective of explicitly outlining how and why a presented text is humorous Hessel et al. (2023).
@@
In addition to the general focus on joke detection, existing works on all humour processing tasks (i.e., detection, explanation, and generation) have mostly focussed on short-form puns such as ”Yesterday I accidentally swallowed some food colouring. The doctor says I’m OK, but I feel like I’ve dyed a little inside” He et al. (2019), where humour arises from the interpretation of the punning word (i.e., ”dyed”) with a phonetically similar word with different meaning (i.e., ”died”) in the case of heterographic puns, or the polysemy of a word for homographic puns Attardo (2008). Whilst puns are commonplace, the humour is primarily self-contained, requiring only a common-sense understanding of the world (e.g., that food colouring ”dyes” things) and an understanding of the semantics of dictionary words (e.g., ”dye”/”die”). On the other hand, much of the humour encountered online, on TV panel shows, and in stand-up comedy routines, is based on contemporary topical knowledge of evolving pop-culture phenomena and news events, where a full appreciation of the humour relies on potentially esoteric knowledge, rather than common-sense Highfield (2015); Laineste (2002), requiring highly complex reasoning and knowledge retrieval. We present examples of each joke type in Figure 1. The homographic pun exploits the dual meaning of ”croaked” as speech and a euphemism for dying; the heterographic pun relies on the phonetic similarity between ”lint” and ”leant”; the non-topical joke plays on the trope of an absentee father (not) returning like a boomerang; and the topical joke refers to the animal welfare organisation PETA’s high euthanasia rates and a reference to the movie Forrest Gump. Each joke type plays on polysemy, phonetics, social constructs, and esoteric knowledge, respectively.
@@
To address the research gap, we perform an assessment of the ability of Large Language Models (LLMs) to explain a range of joke forms in a zero-shot setting, from commonly studied homographic and heterographic puns, to non-topical incongruity-based jokes and topical humour seen on Reddit. As a result, we present the first work on assessing whether state-of-the-art LLMs are equally able to explain different joke formats, from those that require an understanding of basic semantics and phonetics, to those that require esoteric knowledge of pop-culture and news events, therefore assessing the models’ abilities to perform complex reasoning and knowledge retrieval.
@@
We summarise our main contributions below.222We will release all of our code on GitHub. The central dataset has been supplied as supplementary material. In presenting these contributions, we address the primary research questions of (RQ1) whether different joke formats have an impact on the ability of LLMs to explain humour and (RQ2) by extension, whether existing works on computational humour representative of the different joke formats found in everyday life.
@@
We present a novel, balanced dataset of 600 jokes, categorised into traditional puns (homographic/polysemous and heterographic/phonetic), non-topical Reddit jokes (irony, stereotypes, cultural norms, etc.), and topical Reddit jokes (pop culture, news events, etc.). Each joke is paired with a high-quality, succinct, human-authored reference explanation, enabling the first empirical assessment of how joke format affects the ability of LLMs to explain humour.
@@
We use the aforementioned dataset to assess the accuracy and completeness of 4800 joke explanations from 8 state-of-the-art open- and closed-source LLMs of various sizes, including reasoning models (600 jokes * 8 models). We analyse these explanations using human evaluation and automatic metrics, including using an LLM judge.
@@
We qualitatively evaluate the models’ abilities to understand complex humour through LLM-generated explanations of a topical online joke in the form of a case study.
@@ 2 Related Works
Early humour generation efforts relied on simple template-based systems to create puns Valitutti et al. (2013); Ritchie (2005), whilst contemporary works fine-tune deep learning models on extensive datasets of existing jokes Garimella et al. (2020); Weller et al. (2020); He et al. (2019); Petrović and Matthews (2013). More recently, the use of LLMs has become the primary paradigm Chen et al. (2024); Mittal et al. (2022). Furthermore, Sun et al. (2022b) presented CUP, which turned the focus away from standalone joke generation to context-situated generation. Additionally, work has also been performed on humour-adjacent language forms such as tongue-twisters Loakman et al. (2023b); Keh et al. (2023).
@@
Lastly, humour explanation provides an account for why a given text is funny. Consequently, this task requires models to exhibit extensive reasoning capabilities. Whilst explaining standalone puns usually only requires a model to identify the pun word and its alternative interpretations in the given context, more complex jokes require an understanding of common sense, social constructs, and esoteric knowledge to identify key incongruities. Unlike the task of detection, explanation requires models to not only be able to identify humour, but also ”understand” it to provide a textual explanation. Miller et al. (2017) presented a shared task concerning humour, in which subtask 3 pertains to humour interpretation by tasking models with assigning the correct WordNet sense keys to given instances of pun words. Lima Inácio et al. (2023) tackle the humour explanation task indirectly by investigating what specific elements humour classifiers learn. Regarding the production of natural language explanations, Hessel et al. (2023) perform the task of humour explanation on multimodal data from caption contests in The New Yorker, generating explanations as to why a particular best-rated caption is funny.
@@ 3.1 Existing Datasets
Many existing datasets focus on size, rather than annotation depth and consistency. For example, r/Jokes Weller and Seppi (2020) presents 550k jokes taken from Reddit with metadata such as upvote count, whilst SemEval 2017 Task 7 Miller et al. (2017) consists of 3387 simple puns. Whilst Miller et al. (2017) add annotation of the puns’ meanings via WordNet senses, there are no natural language explanations. Sun et al. (2022a) improve upon the SemEval dataset by crowdsourcing textual explanations for a subset, but remains limited to only puns. Furthermore, Hessel et al. (2023) present a dataset of multimodal humour from The New Yorker cartoon caption contest, providing written explanations and Wikipedia links to required world knowledge. Finally, whilst ExplainTheJoke.com333https://explainthejoke.com/ contains detailed joke explanations, the majority are simple puns, in addition to the explanations lacking objectivity, being highly editorialised for entertainment purposes.
@@ 3.2 Our Dataset
In the following section, we outline the process of creating our own dataset with a larger variation of joke types and concise, objective explanations to form the backbone of our humour analysis. Overall, our dataset consists of 600 jokes, containing 150 of each type: homographic puns, heterographic puns, non-topical Reddit humour, and topical Reddit humour. Figure 2 presents the average joke lengths within each subset, whilst Figure 1 presents examples of each subtype (see Appendix A.6 for further examples).
@@ 3.3 Topical Online Humour
To identify jokes that require knowledge of real-world entities and events, we use a subset of r/Jokes Weller and Seppi (2020). Whilst Reddit contains potentially unsavoury material, the presence of strong opinions on topics such as politics and pop culture phenomena results in a rich array of topical humour. Whilst r/Jokes contains a plethora of different topics, not all joke instances rely on contemporary world knowledge. We therefore perform filtering based on upvotes and the presence of named entities (see Appendix A.2 for more details). Of the filtered jokes, we take the top 150 with the most upvotes. We removed jokes that did not require external knowledge, in addition to meta-level humour that relied on knowing that the joke came from Reddit (where this is not explicit).
@@ 3.4 Non-topical Online Humour
We additionally desire to compare the results of topical humour explanation with other common forms of internet humour that do not require esoteric world knowledge. These jokes often present different forms of humour than the simple puns that have been the focus of most existing works on automatic joke explanation, but require less esoteric knowledge than topical jokes. Consequently, we select another 150 top-scoring jokes from the r/Jokes dataset that do not require knowledge beyond what may be considered ”common sense” (e.g., cultural norms). Both the topical and non-topical subsets are checked by another linguist, who is not an author of this work, to ensure their correct categorisation. In effect, the 300 jokes that form the topical and non-topical categories are the top 300 highest-scoring jokes based on upvote-downvote ratio, categorised by their content.
@@ 3.5 Traditional Puns
We source an additional 300 simple pun-based jokes from SemEval 2017 Task 7 Miller et al. (2017)), which has acted as the source data for many works on computational humour. Specifically, we take the first 150 jokes from both the homographic and heterographic training data for Subtask 3 as the basis of the homographic and heterographic joke subsets.
The jokes contained within the topical and non-topical subsets from Reddit may rely on the features of heterographic or homographic puns. However, the interpretation of these puns requires additional real-world knowledge besides the meaning of the pun words. For example, ”New Teslas don’t come with a new car smell. They come with an Elon Musk.”, requires knowledge regarding Elon Musk’s relationship to Tesla.
@@ 3.6 Ground Truth Joke Explanations
As a central contribution, we write ground-truth explanations for all jokes. Due to being written by an author of this work, we perform quality control using 3 native English-speaking computer science students. Each evaluator assessed 200 joke explanations (50 per type), each seeing different jokes.444Participants were given a 30 GBP Amazon voucher for approximately 1.5 hours of evaluation. We evaluate these explanations using the same evaluation criteria used for human evaluation in §5 (with the rubric presented in Appendix A.1), using a 0-5 Likert scale on the criteria of explanation accuracy and completeness. All evaluators scored all human-authored explanations a maximum score of 5 on both criteria or commented on required improvements, which were implemented. Minor explanation post-editing was required for less than 1% of explanations. Further details are in Appendix A.2.
@@ 4.1 Hypotheses
H1. Traditional puns (homographic and heterographic) will be easier for LLMs to explain than high-rated jokes from Reddit, owing to the former’s frequent reliance on the semantics and phonetics of common words.
H2. Heterographic puns will be more difficult for models to explain than homographic puns due to the former’s reliance on phonetic similarity, which is not explicitly encoded in orthographic text.
H3. Topical humour will be more difficult for models to explain than non-topical Reddit humour, due to the former’s reliance on subtle references to contemporary pop culture and events, rather than common sense reasoning and general knowledge.
@@ 4.2 Models
To assess the ability of state-of-the-art LLMs in this task, we select the following suite of open- and closed-source models: Llama 3.1 (8B and 70B) Dubey et al. (2024), Gemini 1.5 (Pro and Flash) Georgiev et al. (2024), GPT-4o (standard and Mini) Achiam et al. (2024). These models represent a range of sizes, using both the flagship large variants and their lightweight counterparts, allowing the investigation of model size on performance in this task. We additionally use variants of DeepSeek-R1 to identify the abilities of reasoning-specific models. Specifically, we use the 8B and 70B variants based upon Llama 3.1 and 3.3, respectively.
Importantly, all of the selected models were trained on data pertaining to events up to at least 2023, with all jokes in the non-topical and topical subsets being posted to Reddit between 2008 and 2019. Therefore, we are fairly assessing the ability of these models to explain jokes that reference events contained within their training data, investigating the models’ ability to correctly associate this knowledge with the abstract references made within topical jokes.
@@ 4.3 Evaluation Criteria
To assess the quality of joke explanations generated by the models, we propose a scoring rubric with two core criteria, accuracy and completeness, each evaluated on a 6-point scale (0-5). Focusing on such criteria allows the application of the scoring rubric to all assessed joke types (such as real-world references in the topical subset constituting completeness, or pun word definitions in traditional puns). The full rubric can be found in Appendix A.1.
Accuracy entails whether or not the joke explanation contains correct material, which does not negatively score content where parts of the joke have not been explained but rather assesses the presence or absence of hallucinations and misunderstandings in a response. On the other hand, Completeness covers whether the fundamental elements of a joke have been addressed. Consequently, a joke may be fully explained (i.e., Completeness = 5), but additional information has been included which is incorrect (Accuracy 3). Conversely, an explanation may identify that a joke is a pun based on a specific word (i.e., Accuracy = 5), but omit explanations of the pun’s meanings or other required information for understanding (i.e., Completeness 2).
Due to the significant amount of time required for rating 4800 explanations ( 40 hours), 1 author (who is a native speaker of English and holds degrees in Linguistics) rated all explanations. However, a subset of 320 explanations (10 jokes * 4 joke types * 8 models) were re-annotated by 2 third-party computer science students (also native English speakers).555Participants were given a 60 GBP Amazon voucher for approximately 3 hours of evaluation. These annotations were used exclusively to verify the primary evaluator’s reliability. The results showed moderate levels of agreement for both accuracy ( = .574) and completeness ( = .553) using Krippendorff’s alpha Krippendorff (2019), as well as high positive Pearson’s correlations Pearson (1895) for accuracy (r = .765) and completeness (r = .795), with , between the scores assigned by the primary human evaluator and third-party evaluators. Overall, 82% of the Accuracy scores and 84% of the Completeness scores from the primary evaluator and the 2 third-party evaluators differed by 1 or 0 points, resulting in an 80.5% agreement between the main evaluator and third-party evaluators when converted to a binary pass/fail (as in 5.1). Across the whole dataset, we additionally leverage the LLM as a Judge paradigm, using Qwen2.5-72B-Instruct, finding agreement for accuracy ( = .565) and completeness ( = .519) to also be substantial, additionally supported by Pearson’s correlations of .641 and .602, both with , respectively. See Appendix A.7 for details on the LLM judge set up, and Appendix A.4 for a deeper analysis of the ratings from the third-party evaluators and LLM judge.
5 Human Evaluation
When comparing model performance, GPT-4o consistently outperforms all others, demonstrating the highest accuracy and completeness scores across joke types. Interestingly, GPT-4o Mini outperforms Gemini Pro in some specific cases, such as in accuracy on homographic jokes, as well as in both accuracy and completeness for non-topical and topical jokes. Additionally, open-source Llama 70B, outperforms most other models except GPT-4o, whilst the R1 models underperform.
5.1 Explanation Success Rate
For an explanation to be useful, it must be both accurate and complete, rather than scoring highly on only one criterion. As a result, in Figure 4 we present a binary categorisation of explanations, treating scores of 4 or above on both criteria as being ”good” quality explanations (i.e., containing only minimal, non-intrusive errors, and/or omitting minor details). This facilitates a clear comparison of the challenges different joke types pose.
As illustrated in Figure 4, the success rate of explanations varies significantly across different joke types. Homographic jokes, where a word has multiple meanings but identical spelling, consistently yield the highest proportion of successful explanations across nearly all models. This suggests that it is easier to navigate the ambiguity of homographs, likely due to language models being aware of the polysemy of the different pun words. In contrast, heterographic jokes, which involve words that sound the same but are spelt differently, pose more difficulty for all models. GPT-4o is the only model that maintains a relatively high success rate in this category. The challenge likely stems from the additional complexity introduced by needing to recognise the phonetic similarity of the pun word whilst being trained on orthographic tokens. Non-topical jokes also show varying levels of performance. The non-topical subset consists of jokes from Reddit, including puns, incongruities and situational irony that cover a wide range of joke forms that have been rated highly by the Reddit userbase, causing the LLMs to struggle more severely than traditional puns. Lastly, topical jokes, which require contextual awareness of specific events, present the most notable challenges. While GPT-4o generates a relatively high proportion of ”good” explanations, other models are substantially less successful. This is likely due to the subtle references to pop culture and real-world entities in the ”topical” subset presenting a significant challenge to current state-of-the-art LLMs, requiring both high levels of reasoning in addition to a form of information retrieval where the information that needs retrieving is not always explicitly referred to.
We see a unique pattern of performance by the Deepseek R1 models. Whilst R1 70B performs well overall, the model’s reasoning capabilities do not result in it outperforming any other tested model. On the other hand, the 8B variant is the worst-performing model. We hypothesise that this is due to the focus on ”reasoning” surrounding mathematical and coding problems, rather than ”reasoning” in the general sense, such as identifying the common-sense incongruities that frequently cause humour, or ”overthinking” the humour in simple puns. Furthermore, due to the reasoning steps taken, it may be the case that R1 8B begins with a moderately good answer (in line with the other models tested), but due to its uncertainty in explaining the jokes, the reasoning steps introduce many opportunities for hallucination, resulting in the smaller 8B model continuing to follow incorrect reasoning before presenting its final answer, therefore becoming progressively worse.
We perform a logistic regression analysis, which further supports our hypotheses. Model size has a strong and significant impact on explanation quality, with larger variants being far more likely to generate ”good” explanations ( = 1.707, p ¡ 0.001). This reinforces our finding that larger models significantly outperform smaller ones. Regarding joke difficulty, the results align with the predicted ordering. Homographic jokes are the easiest to explain ( = 0.583, p ¡ 0.001), while non-topical ( = -0.511, p ¡ 0.001) and topical jokes ( = -0.574, p ¡ 0.001) are significantly harder. This confirms that jokes requiring external context pose a greater challenge to LLMs than those based on simple semantics and phonetics alone. Overall, the model is highly significant (p ¡ 0.001) and explains a meaningful portion of the variance (pseudo R² = 0.138).
6 Automatic Evaluation
In addition to the previously presented human evaluation, we present the results of automatic metrics in Table 1. Whilst existing reference-based automatic metrics are not well designed for this task, we do confirm our originally hypothesised ordering of joke difficulties, with the explanations of homographic/heterographic puns scoring higher than the non-topical joke subset, which in turn scores higher than the topical jokes subset on most metrics.
| Het. | Hom. | Non-Topical | Topical | |
| SacreBLEU | 7.53 | 8.91 | 6.26 | 5.62 |
[…]
| ROUGE-2 | 0.11 | 0.13 | 0.11 | 0.10 |
| ROUGE-L | 0.24 | 0.27 | 0.24 | 0.22 |
[…]
7 Case Study
Example LLM explanations of a randomly selected joke from the topical subset are shown in Figure 5. We omit the R1 reasoning models for space. The joke of interest refers to a 2018 internet phenomenon called the ”Tide Pod Challenge,” where teenagers would dare each other to consume the toxic laundry pods. The central element to the joke’s explanation is the suggestion that this new ”teenage snack food market” increased sales, ergo generating enough revenue for the company to afford advertising at the Super Bowl event. Interestingly, all of the full-size models inferred the reference to this challenge in their explanations. However, the smaller counterparts consistently omitted this information or presented misinterpretations in their explanations. GPT-4o Mini, for example, explains the joke by suggesting that ”teenagers are known for being messy eaters”, which caused increased sales. On the other hand, Llama 8B and Gemini Flash focus on the idea that the humour stems from ”absurdity”, without making any explicit mention of the real-world phenomenon, thereby failing to inform the user of important context. Further discussion of explanation characteristics can be found in Appendix A.6.
8 Discussion
Similarly, H2 concerned the hypothesis that homographic puns (which rely on polysemy) are easier to explain than heterographic puns, which was found to be the case in our tests. We observed that heterographic puns more frequently lead to hallucination, owing to their reliance on phonetic characteristics, with LLMs having only implicit knowledge of how similar different words may ”sound”.
9 Conclusion
Limitations
In this work, we investigated the extent to which joke format and complexity affect the ability of LLMs to generate accurate and complete explanations. However, the nature of topical humour is that it is ever-evolving, with jokes relating to developing news stories and pop culture events continuously appearing. Consequently, our work focuses only on a subset of possible jokes and does not assess the ability of models to explain highly recent jokes, of which there may be only limited knowledge contained within a given model’s training data (as the volume of reporting and discussion on a given topic grows over time).
Due to the jokes being from secondary sources (i.e., not novel jokes written for the purpose of this work), it is likely that the majority of the jokes have been seen before by the LLMs. However, explanations relating to these jokes are much less likely to be present. For the Reddit jokes in particular, the presence of explanations in the training data is reliant on individual Reddit users asking for an explanation within the relevant thread, and someone replying. Whilst this is possible, we sorted the r/Jokes dataset based on upvotes, and intuitively it is less likely that popular jokes require explanations, with Reddit’s userbase already possessing the required knowledge in most cases (hence the jokes receiving a significant amount of upvotes). We manually verified a subset of the jokes and found readily available explanations to be rare.
Ethics Statement
We believe in and firmly adhere to the Code of Conduct in the performance of this work and the methods involved. Whilst the dataset element of our benchmark contains potentially offensive materials (reflecting the often controversial nature of humour found on Reddit), the aim of this resource is to offer explanations of potentially esoteric jokes in order to clarify where particular parties may be perceiving humour from. In this work, we do not create novel humour, nor present the included dataset as appropriate for training humour generation models. Whilst detoxifying and debiasing language models is a growing interest in the NLP community, humans are fundamentally opinionated and biased in ways that others may take offence at. Whilst we agree language models should not perpetuate these biases, the authors of this work believe that, if queried, language models should be able to make use of knowledge pertaining to human beliefs and opinions in offering impartial, objective explanations of phenomena such as humour. We therefore do not view the creation of an explanation for a joke to be equatable with an endorsement of said joke, but rather as an ideologically neutral task. We additionally consciously write our humour explanations in an objective manner, not lending unjust credence to unverified scandals or presenting opinions as facts.
Appendix A Appendix
A.1 Evaluation Rubric

### v2

@@ Abstract
Humour, as a complex language form, is derived from myriad aspects of life. Whilst existing work on computational humour has focussed almost exclusively on short pun-based jokes, we investigate whether the ability of Large Language Models (LLMs) to explain humour depends on the particular form. We compare models' joke explanation abilities from simple puns to complex topical humour that requires esoteric knowledge of real-world entities and events. To this end, we curate a dataset of 600 jokes across 4 joke types and manually write high-quality explanations. These jokes include heterographic and homographic puns, contemporary internet humour, and topical jokes. Using this dataset, we compare the zero-shot abilities of a range of LLMs to accurately and comprehensively explain jokes of different types, identifying key research gaps in the task of humour explanation. We find that none of the tested models (including reasoning models) are capable of reliably generating adequate explanations of all joke types, further highlighting the narrow focus of most existing works on overly simple joke forms.111CONTENT WARNING: Some of the presented jokes may be considered offensive or distasteful to different individuals.
@@ A Dataset & Analysis of LLM Humour Understanding from Traditional Puns to Topical Jokes
Tyler Loakman1, William Thorne1, Chenghua Lin2 1Department of Computer Science, The University of Sheffield, UK 2Department of Computer Science, The University of Manchester, UK tcloakman1@sheffield.ac.uk wthorne1@sheffield.ac.uk chenghua.lin@manchester.ac.uk
@@ 1 Introduction
The perception and understanding of humour is essential to human interaction and entertainment. Consequently, the ability to correctly process jokes is paramount to achieving a human-like understanding of ambiguity and social context in language. However, most existing work on computational humour has narrowly focussed on the simpler task of humour detection Loakman et al. (2023a); Meaney et al. (2021); Miller et al. (2017), which allows models to learn from confounding signals to detect certain joke types Lima Inácio et al. (2023). On the other hand, humour explanation tasks a computational model with the more challenging objective of explicitly outlining how and why a presented text is humorous Hessel et al. (2023).
@@
In addition to the general focus on joke detection, existing works on all humour processing tasks (i.e., detection, explanation, and generation) have mostly focussed on short-form puns such as "Yesterday I accidentally swallowed some food colouring. The doctor says I'm OK, but I feel like I've dyed a little inside" He et al. (2019), where humour arises from the interpretation of the punning word (i.e., "dyed") with a phonetically similar word with different meaning (i.e., "died") in the case of heterographic puns, or the polysemy of a word for homographic puns Attardo (2008). Whilst puns are commonplace, the humour is primarily self-contained, requiring only a common-sense understanding of the world (e.g., that food colouring "dyes" things) and an understanding of the semantics of dictionary words (e.g., dye/die). On the other hand, much of the humour encountered online, on TV panel shows, and in stand-up comedy routines, is based on contemporary topical knowledge of evolving pop-culture phenomena and news events, where a full appreciation of the humour relies on potentially esoteric knowledge, rather than common-sense Highfield (2015); Laineste (2002), requiring highly complex reasoning and knowledge retrieval.
We present examples of each joke type that we investigate in Figure 1. The homographic pun exploits the dual meaning of "croaked" as both a style of speech and a euphemism for dying; the heterographic pun relies on the phonetic similarity between "lint" and "leant"; the non-topical joke plays on the trope of an absentee father (not) returning like a boomerang; and the topical joke refers to the animal welfare organisation PETA's high euthanasia rates and a reference to the movie Forrest Gump. Each joke type plays on polysemy, phonetics, social constructs, and esoteric knowledge, respectively.
@@
To address this research gap, we assess the ability of Large Language Models (LLMs) to explain this range of joke forms in a zero-shot setting. As a result, we present the first work on assessing whether state-of-the-art LLMs are equally able to explain different joke formats, from those that require an understanding of basic semantics and phonetics, to those that require esoteric knowledge of pop culture and news events, therefore assessing the models' abilities to perform complex reasoning and knowledge retrieval.
@@
We summarise our main contributions below. In presenting these contributions, we address the primary research questions of (RQ1) whether different joke formats have an impact on the ability of LLMs to explain humour and by extension, (RQ2) whether findings from existing work on computational humour are generalisable to the different jokes found in everyday life.
@@
We present a novel, balanced dataset of 600 jokes, categorised into puns, non-topical Reddit jokes, and topical Reddit jokes. Each joke is paired with a high-quality, succinct, human-authored reference explanation, enabling the first empirical assessment of how joke format affects the ability of LLMs to explain humour.222We release our dataset and code at: https://github.com/tylerL404/Comparing-Apples-to-Oranges.
@@
We use the aforementioned dataset to assess the accuracy and completeness of 4800 joke explanations across 8 state-of-the-art open- and closed-source LLMs of various sizes, including reasoning models (600 jokes * 8 models). We analyse these explanations using human evaluation and automatic metrics, including using an LLM judge.
@@
We qualitatively evaluate the models' abilities to understand complex humour through LLM-generated explanations of a topical online joke in the form of a case study.
@@ 2 Related Works
Early humour generation efforts relied on simple template-based systems to create puns Valitutti et al. (2013); Ritchie (2005), whilst contemporary works fine-tune deep learning models on extensive datasets of existing jokes Garimella et al. (2020); Weller et al. (2020); He et al. (2019); Petrović and Matthews (2013). More recently, using different prompting strategies with LLMs has become the primary paradigm for generating jokes Chen et al. (2024); Mittal et al. (2022). Furthermore, Sun et al. (2022b) presented CUP, which turned the focus away from standalone joke generation to context-situated generation. Additionally, work has also been performed on humour-adjacent language forms such as tongue-twisters Loakman et al. (2025, 2023b); Keh et al. (2023).
@@
Lastly, humour explanation provides an account for why a given text is funny. Consequently, this task requires models to exhibit extensive reasoning capabilities. Whilst explaining standalone puns usually only requires a model to identify the pun word and its alternative interpretations in the given context, more complex jokes require an understanding of common sense, social constructs, and esoteric knowledge to identify key incongruities. Unlike the task of detection, explanation requires models to not only be able to identify humour, but also "understand" it to provide a textual explanation. Miller et al. (2017) presented a shared task concerning humour, in which subtask 3 pertains to humour interpretation by tasking models with assigning the correct WordNet sense keys to given instances of pun words. Lima Inácio et al. (2023) tackle the humour explanation task indirectly by investigating what specific elements humour classifiers learn. Regarding the production of natural language explanations, Hessel et al. (2023) perform the task of humour explanation on multimodal data from caption contests in The New Yorker, generating explanations as to why a particular best-rated caption is funny. Finally, Wang et al. (2025) investigate the ability of LLMs to explain a wide range of linguistic phenomena that require non-linear reasoning, such as tautologies and sarcasm.
@@ 3.1 Existing Datasets
Many existing datasets focus on size, rather than annotation depth and consistency. For example, r/Jokes Weller and Seppi (2020) presents 550k jokes taken from Reddit with metadata such as upvote count, whilst SemEval 2017 Task 7 Miller et al. (2017) consists of 3387 simple puns. Whilst Miller et al. (2017) add annotation of the puns' meanings via WordNet senses, there are no natural language explanations. Sun et al. (2022a) improve upon the SemEval dataset by crowdsourcing textual explanations for a subset, but the scope remains limited to only puns. Furthermore, Hessel et al. (2023) present a dataset of multimodal humour from The New Yorker cartoon caption contest, providing written explanations and Wikipedia links to required world knowledge. Finally, whilst ExplainTheJoke.com333https://explainthejoke.com/ contains detailed joke explanations, the majority are simple puns and the explanations frequently lack objectivity, being highly editorialised for entertainment purposes.
@@ 3.2 Our Dataset
In the following section, we outline the process of creating our own dataset with a larger range of joke types and concise, objective explanations to form the backbone of our humour analysis. Overall, our dataset contains 600 jokes, consisting of 150 of each type: homographic puns, heterographic puns, non-topical Reddit humour, and topical Reddit humour. Figure 2 presents the average joke lengths within each subset, whilst Figure 1 presents examples of each subtype (see Appendix A.5 for further examples).
@@ 3.3 Topical Online Humour
To identify jokes that require knowledge of real-world entities and events, we use a subset of r/Jokes Weller and Seppi (2020). Whilst Reddit contains potentially unsavoury material, the presence of strong opinions on topics such as politics and pop culture phenomena results in a rich array of topical humour. Whilst r/Jokes contains a plethora of different topics, not all joke instances rely on contemporary world knowledge. We therefore perform filtering based on upvotes and the presence of named entities (see Appendix A.1 for more details). Of the filtered jokes, we take the top 150 with the most upvotes. We removed jokes that did not require external knowledge, in addition to meta-level humour that relied on knowing that the joke came from Reddit (where this is not explicitly mentioned in the joke itself).
@@ 3.4 Non-topical Online Humour
We additionally desire to compare the results of topical humour explanation with other common forms of internet humour that do not require esoteric world knowledge. These jokes often present different forms of humour than the simple puns that have been the focus of most existing works on automatic joke explanation. Consequently, we select another 150 top-scoring jokes from the r/Jokes dataset that do not require knowledge beyond what may be considered "common sense" (e.g., cultural norms). Both the topical and non-topical subsets are checked by another linguist, who is not an author of this work, to ensure their correct categorisation. In effect, the 300 jokes that form the topical and non-topical categories are the top 300 highest-scoring jokes based on upvote-downvote ratio, categorised by their content.
@@ 3.5 Traditional Puns
We source an additional 300 simple pun-based jokes from SemEval 2017 Task 7 Miller et al. (2017), which has acted as the source data for many works on computational humour. Specifically, we take the first 150 jokes from both the homographic and heterographic training data for Subtask 3 as the basis of the homographic and heterographic joke subsets.
The jokes contained within the topical and non-topical subsets from Reddit may rely on the features of heterographic or homographic puns. However, the interpretation of these puns requires additional real-world knowledge besides the meaning of the pun words. For example, "New Teslas don't come with a new car smell. They come with an Elon Musk.", requires knowledge regarding Elon Musk's relationship to Tesla.
@@ 3.6 Ground Truth Joke Explanations
As a central contribution, we write ground-truth explanations for all jokes. Due to being written by an author of this work, we perform quality control using 3 native English-speaking computer science students. Each evaluator assessed 200 joke explanations (50 per type), each seeing different jokes.444Participants were given a 30 GBP Amazon voucher for approximately 1.5 hours of evaluation. We evaluate these explanations using the same evaluation criteria used for human evaluation in §5 (with the rubric presented in Appendix 4.3), using a 0-5 Likert scale on the criteria of explanation accuracy and completeness. All evaluators scored all human-authored explanations a maximum score of 5 on both criteria or commented on required improvements, which were implemented. Minor explanation post-editing was required for less than 1% of explanations. Further details are in Appendix A.1.
@@ 4.1 Hypotheses
H1. Traditional puns (homographic and heterographic) will be easier for LLMs to explain than high-rated jokes from Reddit, owing to the former's frequent reliance on the semantics and phonetics of common words.
H2. Heterographic puns will be more difficult for models to explain than homographic puns due to the former's reliance on phonetic similarity, which is not explicitly encoded in orthographic text.
H3. Topical humour will be more difficult for models to explain than non-topical Reddit humour, due to the former's reliance on subtle references to contemporary pop culture and events, rather than common sense reasoning and general knowledge.
@@ 4.2 Models
To assess the ability of state-of-the-art LLMs in this task, we select the following suite of open- and closed-source models: Llama 3.1 (8B and 70B) Dubey et al. (2024), Gemini 1.5 (Pro and Flash) Georgiev et al. (2024), GPT-4o (standard and Mini) Achiam et al. (2024). These models represent a range of sizes, using both the flagship large variants and their lightweight counterparts, allowing the investigation of model size on performance in this task. We additionally use variants of DeepSeek-R1 to identify the abilities of reasoning-specific models. Specifically, we use the distilled 8B and 70B variants that are based upon Llama 3.1 and 3.3, respectively.
Importantly, all of the selected models were trained on data pertaining to events up to at least 2023, with all jokes in the non-topical and topical subsets being posted to Reddit between 2008 and 2019. Therefore, we are fairly assessing the ability of these models to explain jokes that reference events contained within their training data, investigating the models' ability to correctly associate this knowledge with the abstract references made within topical jokes.
@@ 4.3 Evaluation Criteria
To assess the quality of joke explanations generated by the models, we propose a scoring rubric with two core criteria, accuracy and completeness, each evaluated on a 6-point scale (0-5). Focusing on such criteria allows the application of the scoring rubric to all assessed joke types (such as real-world references in the topical subset constituting completeness, or pun word definitions in traditional puns). The rubric is presented below: